# Fisher-Rao Gradient Flows of Linear Programs and State-Action Natural Policy Gradients

### v1

@@ 5.2.1. Experimental setup and findings
$\theta_{k+1}=\theta_{k}+\eta\cdot\frac{G(\theta_{k})^{+}\nabla R(\theta_{k})}{\lVert G(\theta_{k})^{+}\nabla R(\theta_{k})\rVert},$ (5.20)
where $\eta>0$. This ensures $\lVert\theta_{k+1}-\theta_{k}\rVert=\eta$ and corresponds to using an adaptive step size
$\eta_{k}=\frac{\eta}{\lVert G(\theta_{k})^{+}\nabla R(\theta_{k})\rVert}.$
Consequently, we consider the virtual time $t_{k}\coloneqq\sum_{l=0}^{k-1}\eta_{l}$ in our experiments as $\theta_{k}\approx\tilde{\theta}_{t_{k}}$ if $(\tilde{\theta}_{t})_{t\geq 0}$ solves the natural policy gradient flow. Figure 3 plots the suboptimality gap $R^{\star}-R(\theta_{t})$ as well as the KL-divergence $D_{\operatorname{KL}}(d^{\star},d_{t})$ for the two different natural policy gradient methods and the same $30$ random initializations. Additionally, the gray dashed line indicates the exponential decay rate $O(e^{-\Delta t})=O(e^{-\Delta_{\operatorname{K}}t})$ guaranteed by Corollary 5.8 and by [27], respectively. We see that for all trajectories both the suboptimality gap $R^{\star}-R(\theta_{t})$ as well as the KL-divergence to the optimal state-action distribution $D_{\operatorname{KL}}(d^{\star},d_{t})$ decay at this guaranteed rate for both the state-action and Kakade’s natural policy gradient method.
@@
Figure 3. Shown are the suboptimality gap $R^{\star}-R(\theta_{t})$ (top row) and the KL-divergence $D_{\operatorname{KL}}(d^{\star},d_{t})$ (bottom row) for the state-action NPG (left column) and Kakade’s NPG (right column) plotted in a logarithmic scale, along with the predicted exponential decay $e^{-\Delta t}=e^{-\Delta_{\operatorname{K}}t}$ (dashed line), see Corollary 5.8 and [27] for state-action and Kakade’s NPG, respectively.
@@ 5.2.2. A second example and non-tightness
We complement our computational example by studying the same Markov decision process from Figure 2 but with a modified reward vector given by
@@
As above we can compute the three constants $\delta$, $\Delta$ and $\Delta_{\operatorname{K}}$, and obtain $\delta\approx 0.5326$, $\Delta\approx 0.5789$ and $\Delta_{\operatorname{K}}=1.1$. Here again $\delta<\Delta$ as it is always guaranteed under the 5.2, see Remark 3.4. Further, in this example we have $\Delta<\Delta_{\operatorname{K}}$. We conduct the same experiment as before and report the findings in Figure 4. In the plots concerning the state-action natural policy gradient, we plot both the guaranteed decay $O(e^{-\Delta t})$ (gray dashed line) and the decay $O(e^{-\Delta_{\operatorname{K}}t})$ guaranteed for Kakade’s natural policy gradient (gray dotted line). We see that both methods exhibit the convergence rate $O(e^{-\Delta_{\operatorname{K}}t})$. In particular, this indicates that our convergence analysis of the Fisher-Rao gradient, although improving on known results, is still not tight for general problems.
[…]
Figure 4. Shown are the suboptimality gap $R^{\star}-R(\theta_{t})$ (top) and the KL-divergence $D_{\operatorname{KL}}(d^{\star},d_{t})$ (bottom) for the state-action NPG (left) and Kakade’s NPG (right); shown are also the guaranteed exponential decay rates $e^{-\Delta t}$ for the state-action NPG (dashed line) and $e^{-\Delta_{\operatorname{K}}t}$ for Kakade’s NPG (dotted line). Although the guarantees are different, both methods exhibit the same fast decay rate.
@@ 6. Conclusion and Outlook
We study Fisher-Rao gradient flows of linear programs and show they converge linearly with an exponent that depends on the geometry of the linear program. This yields an estimate on the error introduced by entropic regularization of the linear program, which improves existing guarantees. We extend this analysis to natural gradient flows for general parametrized measure models and show they converge at a sublinear rate $O(\frac{1}{t})$ up to an approximation error and mismatch of the trajectory to the solution $\chi^{2}$-divergence. In particular, our results yield $O(\frac{1}{t})$ convergence of state-action natural policy gradients without regularization under function approximation and linear convergence of state-action natural policy gradients for general tabular parametrizations. Finally, we provide computational examples illustrating our results.
@@ Acknowledgments
The project originated when JM was a PhD student at the International Max Planck Research School Mathematics in the Sciences at MPI MiS with additional support from the Evangelisches Studienwerk Villigst e.V.. GM has been supported in part by NSF CAREER 2145630, NSF 2212520, DFG SPP 2298 project 464109215, ERC 757983, BMBF in DAAD project 57616814.
@@ Appendix A Auxiliary results
For the sake of completeness, we provide the proofs of some auxiliary results.
@@
This is generalization of [59, Lemma 3.6], which covers the case that $F$ consists of a single vertex. We will show that
@@
see [60, Proposition 2.30]. Hence, we can choose convex weights $\lambda_{i}$ such that
@@
See 3.11
Proof.
If $F^{\star}\neq P$, then surely $c^{\top}\mu^{\star}-c^{\top}\mu>0$ for some vertex $\mu$, which implies $\Delta>0$. Now we can proceed analogously to the proof of Lemma 3.8. To simplify notation let us define the set $E\coloneqq\{\mu-\mu^{\star}:\mu\in N(\mu^{\star})\setminus F^{\star},\mu^{\star}\in\operatorname{vert}(F^{\star})\}$ of edges such that exactly one of the two endpoints is contained in $F^{\star}$. Then, the polytope $P$ is contained in
see Lemma A.1 and hence we can write $\mu\in P$ as $\mu=\mu^{\star}+\sum_{e}\alpha_{e}e$ for some $\mu^{\star}\in F^{\star}$. Just like in the proof of Lemma 3.8 we obtain
$\Delta\lVert\mu^{\star}-\mu\rVert_{1}\leq\Delta\sum_{e}\alpha_{e}\lVert e\rVert_{1}\leq\sum_{e}\langle v,-e\rangle=\langle v,\mu^{\star}\rangle-\langle v,\mu\rangle=f^{\star}-\langle v,\mu\rangle.$
@@
The information projection $\hat{\mu}\in F$ is characterized by
$D_{\operatorname{KL}}(\hat{\mu},\mu)=\min_{\mu^{\prime}\in F}D_{\operatorname{KL}}(\mu^{\prime},\mu).$
Assume that $\hat{\mu}\in\partial F$, then $\mu_{x_{0}}=0$ for some $x_{0}\in\mathbb{X}$. Consider now $v\in\mathbb{R}^{\mathbb{X}}$ such that $\hat{\mu}+tv\in\operatorname{int}(F)$ for $t>0$ small enough, then surely $v_{x_{0}}>0$. By convexity of the KL-divergence (or simply of $s\mapsto s\log s$), we have
$D_{\operatorname{KL}}(\hat{\mu},\mu)\geq D_{\operatorname{KL}}(\hat{\mu}+tv,\mu)+t\partial_{t}D_{\operatorname{KL}}(\hat{\mu}+tv,\mu),$
where $\partial_{t}D_{\operatorname{KL}}(\hat{\mu}+tv,\mu)\to-\infty$ for $t\to 0$. Hence, this shows $D_{\operatorname{KL}}(\hat{\mu},\mu)>D_{\operatorname{KL}}(\hat{\mu}+tv,\mu)$ for $t$ small enough contradicting that $\hat{\mu}$ is the information projection of $\mu$. ∎

### v2

@@ 5.2. Computational examples
$\theta_{k+1}=\theta_{k}+\eta\cdot G(\theta_{k})^{+}\nabla R(\theta_{k}),$ (5.16)
with stepsize $\eta>0$, where we choose $\eta=10^{-2}$ in our experiments. Intuitively, we expect $\theta_{k}\approx\tilde{\theta}_{\eta k}$ if $(\tilde{\theta}_{t})_{t\geq 0}$ solves the natural policy gradient flow.
5.2.1. A first example with tightness
Figure 3 plots the suboptimality gap $R^{\star}-R(\theta_{k})$ as well as the KL-divergence $D_{\operatorname{KL}}(d^{\star},d_{\theta_{k}})$ for the two different natural policy gradient methods and the same $30$ random initializations. Additionally, the gray dashed line indicates the exponential decay rate $O(e^{-\Delta\eta k})=O(e^{-\Delta_{\operatorname{K}}\eta k})$ guaranteed by Corollary 5.8 and by [29], respectively. We see that for all trajectories both the suboptimality gap $R^{\star}-R(\theta_{k})$ as well as the KL-divergence to the optimal state-action distribution $D_{\operatorname{KL}}(d^{\star},d_{\theta_{k}})$ decay at this guaranteed rate for both the state-action and Kakade’s natural policy gradient method.
@@ 5.2.1. A first example with tightness
Figure 3. Shown are the suboptimality gap $R^{\star}-R(\theta_{t})$ (top row) and the KL-divergence $D_{\operatorname{KL}}(d^{\star},d_{t})$ (bottom row) for the state-action NPG (left column) and Kakade’s NPG (right column) plotted in a logarithmic scale, along with the predicted exponential decay $e^{-\Delta\eta k}=e^{-\Delta_{\operatorname{K}}\eta k}$ (dashed line), see Corollary 5.8 and [29] for state-action and Kakade’s NPG, respectively.
@@ 5.2.2. A second example and non-tightness
We complement our computational example by studying the same Markov decision process from Figure 2 but changing the reward vector according to $r(s_{1},a_{2})=3$. As above we can compute the three constants $\delta$, $\Delta$ and $\Delta_{\operatorname{K}}$, and obtain $\delta\approx 0.5326$, $\Delta\approx 0.5789$ and $\Delta_{\operatorname{K}}=1.1$. Here again $\delta<\Delta$ as it is always guaranteed under the 5.2, see Remark 3.4. Further, in this example, we have $\Delta<\Delta_{\operatorname{K}}$. We conduct the same experiment as before and report the findings in Figure 4. In the plots concerning the state-action natural policy gradient, we plot both the guaranteed decay $O(e^{-\Delta\eta k})$ (gray dashed line) and the decay $O(e^{-\Delta_{\operatorname{K}}\eta k})$ guaranteed for Kakade’s natural policy gradient (gray dotted line). We see that both methods exhibit the convergence rate $O(e^{-\Delta_{\operatorname{K}}\eta k})$. In particular, this indicates that our convergence analysis of the Fisher-Rao gradient, although improving on known results, is still not tight for general problems.
@@
Figure 4. Shown are the suboptimality $R^{\star}-R(\theta_{k})$ (top) and KL-divergence $D_{\operatorname{KL}}(d^{\star},d_{\theta_{k}})$ (bottom) for the state-action NPG (left) and Kakade’s NPG (right); shown are also the guaranteed exponential decay rates $e^{-\Delta\eta k}$ for the state-action NPG (dashed line) and $e^{-\Delta_{\operatorname{K}}\eta k}$ for Kakade’s NPG (dotted line). Although the guarantees are different, both methods exhibit the same fast decay rate.
@@ 6. Conclusion and Outlook
We study Fisher-Rao gradient flows of linear programs and show they converge linearly with an exponent that depends on the geometry of the linear program. This yields an estimate on the error introduced by entropic regularization of the linear program, which improves existing guarantees. We extend this analysis to natural gradient flows for general parametrized measure models and show they converge at a sublinear rate $O(\frac{1}{t})$ up to an approximation error and mismatch of the trajectory to the solution measure in the $\chi^{2}$-divergence. In particular, our results yield $O(\frac{1}{t})$ convergence of state-action natural policy gradients without regularization under function approximation and linear convergence of state-action natural policy gradients for general tabular parametrizations. Finally, we provide computational examples illustrating our results.
@@ Acknowledgments
The project originated when JM was a PhD student at the International Max Planck Research School Mathematics in the Sciences at MPI MiS with additional support from the Evangelisches Studienwerk Villigst e.V.. SC and JM acknowledge funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under the project number 442047500 through the Collaborative Research Center Sparsity and Singular Structures (SFB 1481). GM has been supported in part by NSF CAREER 2145630, NSF 2212520, DFG SPP 2298 project 464109215, ERC 757983, and BMBF in DAAD project 57616814.
@@ Appendix A Auxiliary results
This is a generalization of [64, Lemma 3.6], which covers the case that $F$ consists of a single vertex. We will show that
@@
see [65, Proposition 2.30]. Hence, we can choose convex weights $\lambda_{i}$ such that
@@
Note that $\hat{\mu}\in F$ is characterized by $D_{\operatorname{KL}}(\hat{\mu},\mu)=\min_{\mu^{\prime}\in F}D_{\operatorname{KL}}(\mu^{\prime},\mu)$. Assume that $\hat{\mu}\in\partial F$, then $\mu_{x_{0}}=0$ for some $x_{0}\in\mathbb{X}$. Consider now $v\in\mathbb{R}^{\mathbb{X}}$ such that $\hat{\mu}+tv\in\operatorname{int}(F)$ for $t>0$ small enough, then surely $v_{x_{0}}>0$. By convexity of the KL-divergence, we have $D_{\operatorname{KL}}(\hat{\mu},\mu)\geq D_{\operatorname{KL}}(\hat{\mu}+tv,\mu)+t\partial_{t}D_{\operatorname{KL}}(\hat{\mu}+tv,\mu)$, where $\partial_{t}D_{\operatorname{KL}}(\hat{\mu}+tv,\mu)\to-\infty$ for $t\to 0$. This shows $D_{\operatorname{KL}}(\hat{\mu},\mu)>D_{\operatorname{KL}}(\hat{\mu}+tv,\mu)$ for $t$ small enough contradicting that $\hat{\mu}$ is the information projection of $\mu$. ∎